### v3

@@ Abstract
Keywords:
Portrait Image Animation Facial Expression Control 3D-aware Synthesis
@@ 1 Introduction
Portrait image animation aims to generate a video of a given source identity with the driving motion. It has received a lot of attention due to the potential of virtual human services, such as cross-lingual film dubbing [31, 13], virtual avatar chatting [41, 70], and video conferencing [58, 56]. In these scenarios, it is essential to transfer the facial expression (e.g., eye-blinking, lip motion, etc.) from different person, i.e., cross-identity transfer, while preserving the source identity. However, it is challenging due to the ambiguity between appearance and expression [18] and the lack of paired data (e.g., different faces with the same expression) for disentanglement representation learning [42].
@@
Most 2D-based methods rely on image warping [52, 71, 65, 59, 24], which warps the source image to the driving image by estimating the motion between them. To impose a bottleneck for the motion representation, they encode the motion into the difference between sparse key-points [52, 71, 24] or latent codes [59], which are trained in an unsupervised manner. However, in this scenario, the facial expressions are encoded into the motion space as well, in terms of local motion, which tends to be neglected due to the relatively large head motions. Furthermore, since the facial expression and the appearance are highly entangled in the image space, cross-identity expression transfer often involves the source appearance change. DPE [42] tackles this entanglement issue by proposing a self-supervised disentanglement learning framework based on cycle-consistency learning [72]. However, it shows temporal inconsistency in the generated video due to its instability of cycle-consistency learning.
@@
Another line of works [34, 38, 67, 35] explores facial expression control in 3D space using the neural radiance fields (NeRFs) [40]. They leverage pre-trained latent representation of 3D GAN [10] for 3D facial prior where they design the expression in terms of latent code [38, 67] or predict deformation field [43] to deform the well-constructed 3D representation, such as tri-plane [34, 38, 67, 35]. However, the latent code cannot faithfully reconstruct the source identity [38], and the point-wise deformation fields to those 3D representations yield video-level artifacts, such as flickers [34].
@@
In this paper, we address the appearance-expression entanglement issue by proposing a contrastive pre-training framework over video datasets that produces appearance-free facial expressions with an orthogonal structure. Armed with this representation, we build a one-shot 3D-aware portrait image animation method, namely Export3D, which controls the facial expression and 3D camera view of a given source image without appearance swap. To achieve this, we design a generator architecture consisting of vision transformer (ViT) and convolution layers [56, 17, 44] that directly generates the tri-planes from the source image and driving expression parameters. Instead of predicting the deformation fields for the expression, we introduce an expression adaptive layer normalization (EAdaLN) which can effectively transfer the driving expression to the source image. The main contributions of this work are summarized as follows:
@@ 2 Related Works
3D-aware image synthesis aims to generate images with explicit camera pose control [50, 10, 21, 15, 11, 61]. This is achieved by conditioning the camera pose parameter into generative features, which are then rendered into an RGB image through differentiable volume rendering [36, 40, 43, 6]. This rendering technique has integrated with adversarial learning [20, 50, 10, 21, 15, 11, 61] to learn 3D view consistency from the unposed dataset. GRAM [15] generates a multi-view consistent image by learning the radiance field on a set of 2D surface manifolds. AniFaceGAN [60] further learns the deformation fields [43] for the facial expression on these manifolds [15] for explicit facial expression control. EG3D [10] introduces a tri-plane representation that provides a strong 3D position encoding with neural volume rendering and become the one of the most prominent representation in this field. However, these methods generate portrait images from noise, requiring further process for real image manipulation.
Relying on the generateive power of EG3D, several works [63, 66, 54, 38, 33, 7, 68, 56, 67] extend 2D GAN-inversion [1, 2, 55, 48] methods, which is challenging due to the multi-view consistency for a single-view image. Specifically, based on facial symmetry, SPI [66] utilizes horizontally flipped images for pseudo supervision to the occluded facial region. However, it requires multi-stage latent code optimizations. GOAE [68] proposes an encoder-based inversion for EG3D which enhances multi-view consistency via an occlusion-aware tri-plane mixing module. Live3DPortrait [56] can reconstruct multi-view consistent portrait images by leveraging the synthetic data of pre-trained EG3D to provide multi-view supervision. However, these methods cannot explicitly manipulate the expression of the source image.
We propose a tri-plane generator architecture that can generate the tri-plane of a given source image with explicit expression control. Inspired by [56, 44], we design this generator with ViT and convolution layers [17], and directly inject expression parameters into the tri-plane generating process through the expression adaptive layer normalization (EAdaLN). By leveraging the strong power of NeRF [40, 10, 56, 67, 54], we decode the generated tri-plane into multi-view images with explicit expression manipulation.
@@ 2.2 Portrait Image Animation
Portrait image animation, or face reenactment, is a task that animates a given source image according to the input driving condition, either audio [45, 31, 41, 70, 37, 22] or image [52, 71, 58, 59, 65]. Specifically, image-driven methods transfer the motion of the driving image into the source image by learning the motion between them. Most works [52, 71, 58] use facial key-points as a pivot representation to be aware of motion via the key-point displacement. FOMM [52] estimates facial key-points in an unsupervised manner, approximating the motion through the first-order Taylor expansion. LIA [59] encodes a motion in terms of latent codes by introducing an orthonormal basis as a motion dictionary. However, the local motion (e.g., facial expression) and the global motion (e.g., head motion) are still entangled in those representations. DPE [42] proposes a bidirectional cyclic training strategy to decouple the pose and expression within the latent codes, while it produces video-level artifacts due to the instability of the cycle-consistency learning.
To explicitly control the facial expression, several works leverage the expression parameters of 3D morphable models (3DMM) [8] in 2D [19, 65] or 3D spaces [38, 35, 34]. StyleHEAT [65] uses 3DMM to warp 2D spatial features of pre-trained StyleGAN2 [28] while yielding texture sticking. OTAvatar [38] proposes a one-shot test-time optimization method that optimizes identity codes of a single source image and learns expression-aware motion latent codes in the latent space of pre-trained EG3D. HiDe-NeRF [34] and NOFA [67] take a different way by predicting an expression-aware deformation field [43] that deforms the tri-plane [10] reconstructed from the source image.
@@ 3 Methods
First of all, we formulate our portrait animation method, Export3D. Given a source image $S\in\mathbb{R}^{3\times H\times W}$, our method transfers the facial expression and camera view of a driving image $D\in\mathbb{R}^{3\times H\times W}$ with the expression and camera parameters, respectively. We employ a tri-plane [10] as the intermediate feature representation, providing a strong 3D position information for differentiable volume rendering [40, 36]. We directly generate an expression-transferred tri-plane from the source image and the driving expression parameter [8, 3] through expression adaptive layer normalization (EAdaLN) (Sec. 3.2). Based on the observation that the expression parameter still contains the appearance information, we propose a pre-training framework using contrastive learning to obtain the appearance-free expression, which forms an orthogonal structure for different expressions (Sec. 3.1). The expression-transferred tri-plane is rendered into a 3D-aware image through the differentiable volume rendering, and then super-resolved into the final output (Sec. 3.3).
@@ 3.1 Contrastive Learned Basis Scaling (CLeBS)
Natural speaking style comes from the the non-verbal component, such as eye blinking. To explicitly control the expression of the generated face, we utilize the expression parameter $\beta\in\mathbb{R}^{64}$ from the widely used 3D morphable models (3DMM) [8] in 3D face reconstruction. However, simply using those parameters for transferring the other person’s expression fails to preserve the facial identity of the source face.
Disentangling Expression and Appearance.  In 3DMM-based face reconstruction, the identity-appearance has been rarely explored. However, [18] shows that a 3D face shape can be reconstructed only using the expression parameters not using the shape parameters, or vice versa. We also observe that the expression parameter of 3DMM is highly entangled with the appearance (Fig. 8(a)), resulting in an undesirable appearance swap when transferring the cross-identity expressions. We assume that the expression parameter needs to be refined to represent pure facial expressions. To address this issue, we propose a contrastive learning based pre-training framework [12, 23, 46, 41] on video dataset to discard the appearance information hidden in the expression parameter. Specifically, given a video sequence $\{X_{i}\}_{i=1}^{N}$ and its corresponding expression sequence $\{\beta_{i}\}_{i=1}^{N}$, we sample an aligned image-expression pair $(X_{k},\beta_{k})$ for the positive and the non-aligned pairs for the negatives as illustrated in Fig. 2. The images and the expressions are mapped into $d$-dimensional representations, and the distance between the positive (or negative) representation pairs is minimized (or maximized) via the following contrastive objective $\mathcal{L}_{cl}$:
@@
where $f_{I}(\cdot)$ is an image encoder, $f_{e}(\cdot)$ is an expression encoder, $\tau$ is the temperature, and $\cos(\cdot,\cdot)$ is the cosine similarity, respectively. Since all samples are from the same video, they share the same appearance, thereby Eq. 1 enforces the encoders to learn appearance-free expression.
Moreover, for designing the expression encoder $f_{e}(\cdot)$, we focus on the orthogonal structure of 3DMM [8] that controls different expressions along different orthogonal directions. To provide the appearance-free expression with the orthogonal structure, we introduce a learned orthonormal basis $V$:
@@
We apply QR-decomposition [59] to a learned weight ($\in\mathbb{R}^{d\times n}$) to explicitly compute the orthonormal basis $V\in\mathbb{R}^{d\times n}$. In this space, an expression is a linear combination of the basis $V=\{\textbf{v}_{i}\}_{i=1}^{n}$ where the coefficient $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ is responsible for the intensity of each expression direction. We call our encoder $f_{e}(\cdot)$ a learned basis scaling (LeBS) module with contrastive pre-training (CLeBS). Once CLeBS is pre-trained with Eq. 1, no further training is required as illustrated in Fig. 1 and Fig. 3, and the image encoder $f_{I}(\cdot)$ is never used after then.
@@ 3.2 Hybrid Tri-plane Generator
where $\text{T}_{ij}\in\mathbb{R}^{32\times\frac{H}{2}\times\frac{W}{2}}$ is the 32-dimensional feature of $\frac{H}{2}\times\frac{W}{2}$ resolution on the $ij$-plane. EG3D [10] utilizes StyleGAN2 [28] to generate the tri-plane from a noise, forming the style latent space $\mathcal{W}\subseteq\mathbb{R}^{512}$. Several works [63, 66, 54, 38, 33, 7, 67] extend the 2D GAN-inversion methods [1, 2, 48, 55, 49] to 3D GAN-inversion in terms of reconstructing the tri-plane from the style latent code.
@@
However, these methods often face challenges in preserving facial identity since the style latent code lacks the capacity for encoding spatial information and person-specific visual details. We directly generate the expression-transferred tri-plane $\text{T}_{\beta_{D}}(S)$ from the source $S$ and the driving expression $\beta_{D}\in\mathbb{R}^{64}$ to reconstruct the driving $D$. Inspired by Live3DPortrait [56], we construct the tri-plane generator with ViT and convolution [17, 62]. Specifically, we convert $S\in\mathbb{R}^{3\times H\times W}$ into a visual feature in $\mathbb{R}^{\frac{h}{4}\times\frac{H}{2^{3}}\times\frac{W}{2^{3}}}$ through a stack of convolutional blocks, and then merge it into the $h$-dimensional $\frac{H\cdot W}{2^{8}}$ visual tokens through a overlap patch merge operator [62]. These tokens and driving expression are processed through a conditional ViT [17, 44, 62] blocks, namely EAdaLN-ViT, where the expression modulates [44] the visual tokens through expression adaptive layer normalization (EAdaLN) as illustrated in Fig. 3. EAdaLN is applied right before the multi-head self-attention (MSA) and the mix feed-forward network (Mix-FFN)[62] of each ViT block to inject the semantic expression into the visual tokens:
@@
where $x$ is the input visual token, $\text{LN}(\cdot)$ is the layer normalization [5], $\sigma(\beta^{\prime}_{D})$ and $\mu(\beta^{\prime}_{D})$ are the $h$-dimensional scale and shift factors computed from $\beta^{\prime}_{D}=f_{e}(\beta_{D})\in\mathbb{R}^{d}$, respectively. To efficiently propagate the visual tokens to the higher resolution, we upsample the visual tokens with pixel shuffle [56] followed by the Gaussian low-pass filter [27]. We experimentally find that the tokens and the pixel shuffle produce grid artifacts, challenging to eliminate in the image space. Employing low-pass filters effectively mitigates these artifacts by smoothing the borderline artifacts over the coordinate. Lastly, we use ViT and convolutional blocks to output the tri-plane $\text{T}_{\beta_{D}}(S)$:
@@
Note that our method does not query the expression parameter to estimate the motion [42, 65, 19], rather it is used as the multi-dimensional label. To stabilize the tri-plane generation, we incorporate the online exponential moving average (EMA) over tri-plane $\text{T}_{EMA}$ which is added to the generated tri-plane. Please refer to supplementary materials for detailed architectures.

### v4

@@ 1 Introduction
Portrait image animation aims to generate a video of a given source identity with the driving motion. It has received a lot of attention due to the potential of virtual human services, such as cross-lingual film dubbing [stylelipsync, videoretalking], virtual avatar chatting [styletalker, sadtalker], and video conferencing [osfv, live3d]. In these scenarios, it is essential to transfer the facial expression (e.g., eye-blinking, lip motion, etc.) from different person, i.e., cross-identity transfer, while preserving the source identity. However, it is challenging due to the ambiguity between appearance and expression [ambiguity] and the lack of paired data (e.g., different faces with the same expression) for disentanglement representation learning [dpe].
@@
Most 2D-based methods rely on image warping [fomm, tpsmm, styleheat, lia, mcnet], which warps the source image to the driving image by estimating the motion between them. To impose a bottleneck for the motion representation, they encode the motion into the difference between sparse key-points [fomm, tpsmm, mcnet] or latent codes [lia], which are trained in an unsupervised manner. However, in this scenario, the facial expressions are encoded into the motion space as well, in terms of local motion, which tends to be neglected due to the relatively large head motions. Furthermore, since the facial expression and the appearance are highly entangled in the image space, cross-identity expression transfer often involves the source appearance change. DPE [dpe] tackles this entanglement issue by proposing a self-supervised disentanglement learning framework based on cycle-consistency learning [cyclegan]. However, it shows temporal inconsistency in the generated video due to its instability of cycle-consistency learning.
@@
Another line of works [hidenerf, otavatar, nofa, goavatar] explores facial expression control in 3D space using the neural radiance fields (NeRFs) [nerf]. They leverage pre-trained latent representation of 3D GAN [eg3d] for 3D facial prior where they design the expression in terms of latent code [otavatar, nofa] or predict deformation field [nerfies] to deform the well-constructed 3D representation, such as tri-plane [hidenerf, otavatar, nofa, goavatar]. However, the latent code cannot faithfully reconstruct the source identity [otavatar], and the point-wise deformation fields to those 3D representations yield video-level artifacts, such as flickers [hidenerf].
@@
In this paper, we address the appearance-expression entanglement issue by proposing a contrastive pre-training framework over video datasets that produces appearance-free facial expressions with an orthogonal structure. Armed with this representation, we build a one-shot 3D-aware portrait image animation method, namely Export3D, which controls the facial expression and 3D camera view of a given source image without appearance swap. To achieve this, we design a generator architecture consisting of vision transformer (ViT) and convolution layers [live3d, vit, dit] that directly generates the tri-planes from the source image and driving expression parameters. Instead of predicting the deformation fields for the expression, we introduce an expression adaptive layer normalization (EAdaLN) which can effectively transfer the driving expression to the source image. The main contributions of this work are summarized as follows:
@@ 2 Related Works
3D-aware image synthesis aims to generate images with explicit camera pose control [graf, eg3d, stylenerf, gram, pigan, gramhd]. This is achieved by conditioning the camera pose parameter into generative features, which are then rendered into an RGB image through differentiable volume rendering [neuralvolumes, nerf, nerfies, mipnerf]. This rendering technique has integrated with adversarial learning [gan, graf, eg3d, stylenerf, gram, pigan, gramhd] to learn 3D view consistency from the unposed dataset. GRAM [gram] generates a multi-view consistent image by learning the radiance field on a set of 2D surface manifolds. AniFaceGAN [anifacegan] further learns the deformation fields [nerfies] for the facial expression on these manifolds [gram] for explicit facial expression control. EG3D [eg3d] introduces a tri-plane representation that provides a strong 3D position encoding with neural volume rendering and become the one of the most prominent representation in this field. However, these methods generate portrait images from noise, requiring further process for real image manipulation.
Relying on the generateive power of EG3D, several works [hfgi3d, spi, next3d, otavatar, 3dpose, triplanenet, goae, live3d, nofa] extend 2D GAN-inversion [im2st, im2st++, e4e, psp] methods, which is challenging due to the multi-view consistency for a single-view image. Specifically, based on facial symmetry, SPI [spi] utilizes horizontally flipped images for pseudo supervision to the occluded facial region. However, it requires multi-stage latent code optimizations. GOAE [goae] proposes an encoder-based inversion for EG3D which enhances multi-view consistency via an occlusion-aware tri-plane mixing module. Live3DPortrait [live3d] can reconstruct multi-view consistent portrait images by leveraging the synthetic data of pre-trained EG3D to provide multi-view supervision. However, these methods cannot explicitly manipulate the expression of the source image.
We propose a tri-plane generator architecture that can generate the tri-plane of a given source image with explicit expression control. Inspired by [live3d, dit], we design this generator with ViT and convolution layers [vit], and directly inject expression parameters into the tri-plane generating process through the expression adaptive layer normalization (EAdaLN). By leveraging the strong power of NeRF [nerf, eg3d, live3d, nofa, next3d], we decode the generated tri-plane into multi-view images with explicit expression manipulation.
@@ 2.2 Portrait Image Animation
Portrait image animation, or face reenactment, is a task that animates a given source image according to the input driving condition, either audio [wav2lip, stylelipsync, styletalker, sadtalker, styletalk, adnerf] or image [fomm, tpsmm, osfv, lia, styleheat]. Specifically, image-driven methods transfer the motion of the driving image into the source image by learning the motion between them. Most works [fomm, tpsmm, osfv] use facial key-points as a pivot representation to be aware of motion via the key-point displacement. FOMM [fomm] estimates facial key-points in an unsupervised manner, approximating the motion through the first-order Taylor expansion. LIA [lia] encodes a motion in terms of latent codes by introducing an orthonormal basis as a motion dictionary. However, the local motion (e.g., facial expression) and the global motion (e.g., head motion) are still entangled in those representations. DPE [dpe] proposes a bidirectional cyclic training strategy to decouple the pose and expression within the latent codes, while it produces video-level artifacts due to the instability of the cycle-consistency learning.
To explicitly control the facial expression, several works leverage the expression parameters of 3D morphable models (3DMM) [3dmm] in 2D [pechead, styleheat] or 3D spaces [otavatar, goavatar, hidenerf]. StyleHEAT [styleheat] uses 3DMM to warp 2D spatial features of pre-trained StyleGAN2 [stylegan2] while yielding texture sticking. OTAvatar [otavatar] proposes a one-shot test-time optimization method that optimizes identity codes of a single source image and learns expression-aware motion latent codes in the latent space of pre-trained EG3D. HiDe-NeRF [hidenerf] and NOFA [nofa] take a different way by predicting an expression-aware deformation field [nerfies] that deforms the tri-plane [eg3d] reconstructed from the source image.
@@ 3 Methods
First of all, we formulate our portrait animation method, Export3D. Given a source image $S\in\mathbb{R}^{3\times H\times W}$, our method transfers the facial expression and camera view of a driving image $D\in\mathbb{R}^{3\times H\times W}$ with the expression and camera parameters, respectively. We employ a tri-plane [eg3d] as the intermediate feature representation, providing a strong 3D position information for differentiable volume rendering [nerf, neuralvolumes]. We directly generate an expression-transferred tri-plane from the source image and the driving expression parameter [3dmm, expression] through expression adaptive layer normalization (EAdaLN) (Sec.˜3.2). Based on the observation that the expression parameter still contains the appearance information, we propose a pre-training framework using contrastive learning to obtain the appearance-free expression, which forms an orthogonal structure for different expressions (Sec.˜3.1). The expression-transferred tri-plane is rendered into a 3D-aware image through the differentiable volume rendering, and then super-resolved into the final output (Sec.˜3.3).
@@ 3.1 Contrastive Learned Basis Scaling (CLeBS)
Natural speaking style comes from the the non-verbal component, such as eye blinking. To explicitly control the expression of the generated face, we utilize the expression parameter $\beta\in\mathbb{R}^{64}$ from the widely used 3D morphable models (3DMM) [3dmm] in 3D face reconstruction. However, simply using those parameters for transferring the other person’s expression fails to preserve the facial identity of the source face.
Disentangling Expression and Appearance.  In 3DMM-based face reconstruction, the identity-appearance has been rarely explored. However, [ambiguity] shows that a 3D face shape can be reconstructed only using the expression parameters not using the shape parameters, or vice versa. We also observe that the expression parameter of 3DMM is highly entangled with the appearance (Fig. 8(a)), resulting in an undesirable appearance swap when transferring the cross-identity expressions. We assume that the expression parameter needs to be refined to represent pure facial expressions. To address this issue, we propose a contrastive learning based pre-training framework [simclr, moco, clip, styletalker] on video dataset to discard the appearance information hidden in the expression parameter. Specifically, given a video sequence $\{X_{i}\}_{i=1}^{N}$ and its corresponding expression sequence $\{\beta_{i}\}_{i=1}^{N}$, we sample an aligned image-expression pair $(X_{k},\beta_{k})$ for the positive and the non-aligned pairs for the negatives as illustrated in Fig.˜2. The images and the expressions are mapped into $d$-dimensional representations, and the distance between the positive (or negative) representation pairs is minimized (or maximized) via the following contrastive objective $\mathcal{L}_{cl}$:
@@
where $f_{I}(\cdot)$ is an image encoder, $f_{e}(\cdot)$ is an expression encoder, $\tau$ is the temperature, and $\cos(\cdot,\cdot)$ is the cosine similarity, respectively. Since all samples are from the same video, they share the same appearance, thereby Eq.˜1 enforces the encoders to learn appearance-free expression.
Moreover, for designing the expression encoder $f_{e}(\cdot)$, we focus on the orthogonal structure of 3DMM [3dmm] that controls different expressions along different orthogonal directions. To provide the appearance-free expression with the orthogonal structure, we introduce a learned orthonormal basis $V$:
@@
We apply QR-decomposition [lia] to a learned weight ($\in\mathbb{R}^{d\times n}$) to explicitly compute the orthonormal basis $V\in\mathbb{R}^{d\times n}$. In this space, an expression is a linear combination of the basis $V=\{\textbf{v}_{i}\}_{i=1}^{n}$ where the coefficient $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ is responsible for the intensity of each expression direction. We call our encoder $f_{e}(\cdot)$ a learned basis scaling (LeBS) module with contrastive pre-training (CLeBS). Once CLeBS is pre-trained with Eq.˜1, no further training is required as illustrated in Fig.˜1 and Fig.˜3, and the image encoder $f_{I}(\cdot)$ is never used after then.
@@ 3.2 Hybrid Tri-plane Generator
where $\text{T}_{ij}\in\mathbb{R}^{32\times\frac{H}{2}\times\frac{W}{2}}$ is the 32-dimensional feature of $\frac{H}{2}\times\frac{W}{2}$ resolution on the $ij$-plane. EG3D [eg3d] utilizes StyleGAN2 [stylegan2] to generate the tri-plane from a noise, forming the style latent space $\mathcal{W}\subseteq\mathbb{R}^{512}$. Several works [hfgi3d, spi, next3d, otavatar, 3dpose, triplanenet, nofa] extend the 2D GAN-inversion methods [im2st, im2st++, psp, e4e, pti] to 3D GAN-inversion in terms of reconstructing the tri-plane from the style latent code.
@@
However, these methods often face challenges in preserving facial identity since the style latent code lacks the capacity for encoding spatial information and person-specific visual details. We directly generate the expression-transferred tri-plane $\text{T}_{\beta_{D}}(S)$ from the source $S$ and the driving expression $\beta_{D}\in\mathbb{R}^{64}$ to reconstruct the driving $D$. Inspired by Live3DPortrait [live3d], we construct the tri-plane generator with ViT and convolution [vit, segformer]. Specifically, we convert $S\in\mathbb{R}^{3\times H\times W}$ into a visual feature in $\mathbb{R}^{\frac{h}{4}\times\frac{H}{2^{3}}\times\frac{W}{2^{3}}}$ through a stack of convolutional blocks, and then merge it into the $h$-dimensional $\frac{H\cdot W}{2^{8}}$ visual tokens through a overlap patch merge operator [segformer]. These tokens and driving expression are processed through a conditional ViT [vit, dit, segformer] blocks, namely EAdaLN-ViT, where the expression modulates [dit] the visual tokens through expression adaptive layer normalization (EAdaLN) as illustrated in Fig.˜3. EAdaLN is applied right before the multi-head self-attention (MSA) and the mix feed-forward network (Mix-FFN)[segformer] of each ViT block to inject the semantic expression into the visual tokens:
@@
where $x$ is the input visual token, $\text{LN}(\cdot)$ is the layer normalization [layernorm], $\sigma(\beta^{\prime}_{D})$ and $\mu(\beta^{\prime}_{D})$ are the $h$-dimensional scale and shift factors computed from $\beta^{\prime}_{D}=f_{e}(\beta_{D})\in\mathbb{R}^{d}$, respectively. To efficiently propagate the visual tokens to the higher resolution, we upsample the visual tokens with pixel shuffle [live3d] followed by the Gaussian low-pass filter [stylegan3]. We experimentally find that the tokens and the pixel shuffle produce grid artifacts, challenging to eliminate in the image space. Employing low-pass filters effectively mitigates these artifacts by smoothing the borderline artifacts over the coordinate. Lastly, we use ViT and convolutional blocks to output the tri-plane $\text{T}_{\beta_{D}}(S)$:
@@
Note that our method does not query the expression parameter to estimate the motion [dpe, styleheat, pechead], rather it is used as the multi-dimensional label. To stabilize the tri-plane generation, we incorporate the online exponential moving average (EMA) over tri-plane $\text{T}_{EMA}$ which is added to the generated tri-plane. Please refer to supplementary materials for detailed architectures.